years. a) How many degrees of freedom does the t-statistic have? b) How many degrees of freedom would the t-statistic have if the sample size had been 225? A survey of 25 randomly selected customers found the ages shown (in years). The mean is 33.68 years and the standard deviation is 10.5731 27 49 33 48 49 34 35 37 28 21 26 48 38 35 20 27 18 26 10 38 36 44 49 35 a) The t-statistic has (Simplify your answer.) degrees of freedom. b) If the sample size had been 225, the t-statistic would havedegrees of freedom (Simplify your answer.) 0Please answer a and b
Answer #1

Answer :

Given that :

randomly selected customers = 25

mean of years = 33.68

standard deviation =10.57

sample size has been = 225

(A). now we need to find out degrees of freedom for the t-statistic

when n = 25

we know that the

the degree of freedom = n - 1

= 25 - 1

= 24

\therefore  the degree of freedom for the t-statistic is 24

(B).  now we need to find out degrees of freedom for the t-statistic

when n = 225

we know that the

the degree of freedom = n - 1

= 225 - 1

= 224

\therefore  the degree of freedom for the t-statistic is 224